# Stronger-MAS: Multi-Agent Reinforcement Learning for Collaborative LLMs

### v4

@@ Appendix A Extended Comparison of RL Training for LLM-based MAS
MAPoRL park2025maporl R P
MARFT Liao2025MARFT R S
MARTI zhang2025marti R S+P
CURE wang2025cure S P
@@
Note. For MARFT [Liao2025MARFT], we report the characteristics of version v3 of the framework, corresponding to the preprint available prior to the completion of this work.
As summarized in Tab. 5, StrongerMAS is the only RL-based LLM multi-agent training framework that simultaneously supports both shared and role-specific policies, hybrid sequential–parallel execution, multi-turn interaction, and heterogeneous agent roles, while being validated on multiple task domains and implemented as a general-purpose MAS RL algorithm. In contrast, MAPoRL [park2025maporl] and CURE [wang2025cure] restrict policies to a purely role-shared setting and are tailored to specific tasks, MARFT Liao2025MARFT restricts agents to single-turn sequential interactions, and MARTI zhang2025marti merely introduces basic single-agent RL algorithms (e.g., GRPO) to the MAS setting. . This combination of flexible workflow expressivity (S+R, S+P, heterogeneous roles) and broad, cross-domain evaluation makes StrongerMAS a more faithful and scalable abstraction for training cooperative LLM-based MAS.
@@ B.7 Theoretical Justification for Greedy Turn-Level Transitions
In this section, we formally justify the optimality of greedy selection based on environment-verified rewards. Consider the underlying MDP with optimal action-value function $Q^{*}(s,a)$. The Bellman optimality principle implies that any policy $\pi^{*}$ satisfying $\pi^{*}(s)\in\arg\max_{a}Q^{*}(s,a)$ is optimal [sutton2018rl].
@@ Appendix F Multi turn single agent
Tab. 7 and Tab. 8 report single-agent ablations on Code and Math. For both 1.7B and 8B models, introducing a multi-turn SA variant (i.e., letting one agent repeatedly revise its own answer) brings no consistent benefit over the standard single-turn SA baseline and often degrades performance, which is align with the obeservation in chen2025setsleveragingselfverificationselfcorrection . For example, at the SFT stage on Qwen3-1.7B, LiveCodeBench drops from $11.6$ to $10.4$ and AIME24 from $13.4$ to $3.3$ when switching from single-turn SA to multi-turn SA, with similar trends on AIME25 and Olympiad. After RL, the single-turn SA policy still outperforms its multi-turn counterpart across most Code and Math benchmarks for both model scales. These results support our claim in the main text: in the absence of additional environmental signals or feedback from complementary roles, multi-turn SA interaction is a contrived use of extra turns that departs from the QA-style pretraining regime and fails to translate into improved task performance, in contrast to our multi-agent workflows where multi-turn interaction with structured cross-agent feedback yields clear gains.

### v5

@@ Appendix A Extended Comparison of RL Training for LLM-based MAS
MAPoRL Park et al. [2025a] R P
MARFT Liao et al. [2025] R S
MARTI Zhang et al. [2025] R S+P
CURE Wang et al. [2025a] S P
@@
Note. For MARFT [Liao et al., 2025], we report the characteristics of version v3 of the framework, corresponding to the preprint available prior to the completion of this work.
As summarized in Tab. 5, StrongerMAS is the only RL-based LLM multi-agent training framework that simultaneously supports both shared and role-specific policies, hybrid sequential–parallel execution, multi-turn interaction, and heterogeneous agent roles, while being validated on multiple task domains and implemented as a general-purpose MAS RL algorithm. In contrast, MAPoRL [Park et al., 2025a] and CURE [Wang et al., 2025a] restrict policies to a purely role-shared setting and are tailored to specific tasks, MARFT Liao et al. [2025] restricts agents to single-turn sequential interactions, and MARTI Zhang et al. [2025] merely introduces basic single-agent RL algorithms (e.g., GRPO) to the MAS setting. . This combination of flexible workflow expressivity (S+R, S+P, heterogeneous roles) and broad, cross-domain evaluation makes StrongerMAS a more faithful and scalable abstraction for training cooperative LLM-based MAS.
@@ B.7 Theoretical Justification for Greedy Turn-Level Transitions
In this section, we formally justify the optimality of greedy selection based on environment-verified rewards. Consider the underlying MDP with optimal action-value function $Q^{*}(s,a)$. The Bellman optimality principle implies that any policy $\pi^{*}$ satisfying $\pi^{*}(s)\in\arg\max_{a}Q^{*}(s,a)$ is optimal [Sutton and Barto, 2018].
@@ Appendix F Multi turn single agent
Tab. 7 and Tab. 8 report single-agent ablations on Code and Math. For both 1.7B and 8B models, introducing a multi-turn SA variant (i.e., letting one agent repeatedly revise its own answer) brings no consistent benefit over the standard single-turn SA baseline and often degrades performance, which is align with the obeservation in Chen et al. [2025b] . For example, at the SFT stage on Qwen3-1.7B, LiveCodeBench drops from $11.6$ to $10.4$ and AIME24 from $13.4$ to $3.3$ when switching from single-turn SA to multi-turn SA, with similar trends on AIME25 and Olympiad. After RL, the single-turn SA policy still outperforms its multi-turn counterpart across most Code and Math benchmarks for both model scales. These results support our claim in the main text: in the absence of additional environmental signals or feedback from complementary roles, multi-turn SA interaction is a contrived use of extra turns that departs from the QA-style pretraining regime and fails to translate into improved task performance, in contrast to our multi-agent workflows where multi-turn interaction with structured cross-agent feedback yields clear gains.